A car with mass = 1000 kg is traveling at a velocity of 100 m/s. What’s the car’s momentum?



1
Expert's answer
2021-01-11T11:40:47-0500

By the definition of the momentum, we have:


"p=mv=1000\ kg\cdot 100\ \dfrac{m}{s}=10\cdot10^4\ \dfrac{kgm}{s}."

Answer:

"p=10\cdot10^4\ \dfrac{kgm}{s}."